# Audience Impressions of Narrative Structures and Personal Language Style in Science Communication on Social Media

### v1

@@ 1. Introduction
On social media, science communication strategies emphasize making science engaging and understandable to non-technical audiences and adapt common practices in science communication literature for a social media audience (Gilbert, 2020). A study of science communication on Twitter found these strategies: 1) use of a relatable example – such as using the iridescence of bubbles to motivate the topic “Thin Film Interference” in physics, 2) use of a step-by-step walkthrough that uses 10-12 tweets that uses example to explain every step, and 3) use of personal language – the writer talking about themselves and their experience, often in an personal and conversational way, “Crazy, right? WTH was I thinking ?!?” (Gero et al., 2021a).
But previous work has not evaluated whether these science communication techniques on social media (example, walkthrough, and personal language) are actually helpful in supporting everyday people learn about science. As such, we conduct a study of non-technical readers’ rating of STEM explanations with and without each of these three social media explanation techniques to understand how each factor impacts making science explanations engaging and understandable. To facilitate fair comparisons between each condition, we use an LLM to generate science explanations for 15 topics: 5 STEM fields, each with 3 topics at different levels of complexity. Readers evaluated them for understanding and engagement. Overall, readers preferred explanations with examples over no examples, and personal language over no personal language. However, readers were split between narratives with and without walkthroughs; certain topics were more suitable for narratives without a walkthrough. Even for techniques they did prefer, readers’ personal experience complicate why some preferred narratives without a given technique. There is no one-size-fits-all approach to science communication on social media.
@@ 2.1.2. Walkthrough (W)
Tweetorial structures often use a narrative and signposting to establish a narrative structure. We define these two attributes as the walkthrough (W) technique. Narrative is defined by a series of connected events to explain a given topic. Tweetorials signpost by using transition words like “Firstly” and “Secondly,” or by using a list of questions to help frame the sequence of the Tweetorial. In a Tweetorial about selectivity metrics in college rankings, the author uses the second tweet to list out 3 driving questions for the explanation adn to establish the structure of the Tweetorial: “1. Does “selectivity” actually tell you anything useful about how good your education will be? 2. What does “selectivity” actually measure that is of value to a student? 3. Why do I have the feeling somebody chose this metric cause they just needed more stuff to rank by?” [^2] http://language-play.com/tech-tweets/tweetorial/14
@@ 3.1. HCI for Science Communication
Science communication on social media is informal and often told from a personal perspective to engage a broader audience (Gero et al., 2021b). Previous work has compiled recommendations for effective science communication on social media based on how likely using these strategies will increase engagement: comments, likes, and shares (zhang2022no; Bik et al., 2015; Fontaine et al., 2019). But these works do not consider whether these techniques are actually effective for readers to understand the science. Our work evaluates whether these different science communication strategies on Twitter are effective in helping readers engage with and understand the science.
@@ 4.2.2. Generating Tweetorials with and without Personal Language (H3)
Figure 3 shows an shows a side-by-side annotated example of the experimental condition (EWP) and baseline condition (EW) for the topic of Walker’s Action Decrement Theory. The lack of personal language highlights in yellow in the baseline condition shows the effect of removing personal language from EWP.

### v2

@@ 1. Introduction
On social media, science communication strategies emphasize making science engaging and understandable to non-technical audiences and adapt common practices in science communication literature for a social media audience (Gilbert, 2020). A study of science communication on Twitter found these strategies: 1) use of a relatable example – such as using the iridescence of bubbles to motivate the topic “Thin Film Interference” in physics, 2) use of a step-by-step walkthrough that uses 10-12 tweets which use an example to explain every step, and 3) use of personal language – the writer talking about themselves and their experience, often in a personal and conversational way, “Crazy, right? WTH was I thinking ?!?” (Gero et al., 2021a).
But previous work has not evaluated whether these science communication techniques on social media (example, walkthrough, and personal language) are actually helpful in helping everyday people learn about science. As such, we conduct a study of non-technical readers’ rating of STEM explanations with and without each of these three social media explanation techniques to understand how each factor impacts making science explanations engaging and understandable. To facilitate fair comparisons between each condition, we use an LLM to generate science explanations for 15 topics: 5 STEM fields, each with 3 topics at different levels of complexity. Readers evaluated them for understanding and engagement. Overall, readers preferred explanations with examples over no examples and personal language over no personal language. However, readers were split between narratives with and without walkthroughs; certain topics were more suitable for narratives without a walkthrough. Even for techniques they did prefer, readers’ personal experience complicate why some preferred narratives without a given technique. There is no one-size-fits-all approach to science communication on social media.
@@ 2.1.2. Walkthrough (W)
Tweetorial structures often use a narrative and signposting to establish a narrative structure. We define these two attributes as the walkthrough (W) technique. Narrative is defined by a series of connected events to explain a given topic. Tweetorials signpost by using transition words like “Firstly” and “Secondly,” or by using a list of questions to help frame the sequence of the Tweetorial. In a Tweetorial about selectivity metrics in college rankings, the author uses the second tweet to list out 3 driving questions for the explanation and to establish the structure of the Tweetorial: “1. Does “selectivity” actually tell you anything useful about how good your education will be? 2. What does “selectivity” actually measure that is of value to a student? 3. Why do I have the feeling somebody chose this metric cause they just needed more stuff to rank by?” [^2] http://language-play.com/tech-tweets/tweetorial/14
@@ 3.1. HCI for Science Communication
Science communication on social media is informal and often told from a personal perspective to engage a broader audience (Gero et al., 2021b). Previous work has compiled recommendations for effective science communication on social media based on how likely using these strategies will increase engagement: comments, likes, and shares (Zhang and Lu, 2022; Bik et al., 2015; Fontaine et al., 2019). But these works do not consider whether these techniques are actually effective for readers to understand the science. Our work evaluates whether these different science communication strategies on Twitter are effective in helping readers engage with and understand the science.
@@ 4.2.2. Generating Tweetorials with and without Personal Language (H3)
Figure 3 shows a side-by-side annotated example of the experimental condition (EWP) and baseline condition (EW) for the topic of Walker’s Action Decrement Theory. The lack of personal language highlights in yellow in the baseline condition shows the effect of removing personal language from EWP.